OpenStudy (anonymous):

Another problem that had got me stupd. Hope you can understand it Using a deck of cards and picking 12 out to work with. 1. Hearts Ace 2. Diamond King 3. Club 2 4. Heart 7 5. Spade Queen 6. Spade 4 7. Diamond 4 8. Club 8 9. Heart Kind 10. Spade 2 11. Heart 5 12. Club 9 Questions: A. Find expermental probability of picking a face card B. Find experimental probability of picking a heart C.What theoretical probabililty of picking a face card, replacing it, and then picking another face card DWhat is the theorical probabiliity of picking a hear, not replacing it, then picki

OpenStudy (rajee_sam):

Out of the deck you pick out specifically these 12 cards to work with. In the 12 cards A. what is the P (Face Card)? = 4/12 = 1/3 B what is the P( Heart) = 4/12 = 1/3 C. Probability of (Face Card replacing and Face) = (16/52) * (16/52) = 16/169 D. What is P( Heart not replacing Heart) = (13/52) * (12/51) = 1/17

OpenStudy (rajee_sam):

Theoretical probability will be with the whole deck. Your experiment is just among the 12 cards you picked to work with

OpenStudy (anonymous):

Oh, that explains it. So on C we are doing a whole deck and not just the 12.

OpenStudy (anonymous):

How did you get D. I don't understand

OpenStudy (rajee_sam):

D there are 13 hearts cards so P of getting one is 13/52 for first draw. Second draw you are not replacing the first one. So only 12 are left and only 51 total cards left. So picking a heart now will have a prob. = 12/51. The combined probability will be 13 * 12 / 52 * 51 = you simplify you get 1/17

OpenStudy (anonymous):

I got it. Now I have to still write. the diffenence between c and D to show dependent events and independent events. Any ideas. then write a compund problem using dependent events involving standard deck of cards and then solve along with a independent problem.

OpenStudy (anonymous):

Thank you for your help. Have a good day!